OpenStudy (kaylala):

proving identities: (tan y + sec^2 y - tan^2 y)/sec y = sin y + cos y

OpenStudy (anonymous):

(tan y + sec^2 y - tan^2 y)/sec y 1+tan^2 y=sec^2 y =(tan y + 1+tan^2 y - tan^2 y)/sec y =(tan y + 1)/sec y=tan y/sec y+1/sec y sec y=1/cos y , tan y =sin y/cos y =sin y + cos y

OpenStudy (anonymous):

From L.H.S (Tan Y + Sec² Y - Tan²Y) /Sec Y = (Sin/Cos Y + 1) / 1/Cos Y (As we know Tan Y = Sin/Cos Y = Cos Y* Sin/Cos Y + Cos Y *1 Sec²-Tan²=1 = Sin Y + Cos Y = R.H.S Sec Y = 1/Cos Y )